0.8 Six mass spring systems are presented in the table below. 27 27 27 F927 a pairs of systems, if any, are dynamically equivalent lic., share the same wo]?
Answer #1

wo = sqrt (k/M)

Therefore,

we can say that pair of systems (A, E) and ( B, F) are dynamically equivalent systems.

for pair (A, E), wo = sqrt (3/9) = sqrt (9/27) = 0.57735

for pair (B, F), wo = sqrt (9/3) = sqrt (27/9) = 1.732